# Global graph features unveiled by unsupervised geometric deep learning

### v2

@@ Abstract
Graphs provide a powerful framework for modeling complex systems, but their structural variability poses significant challenges for analysis and classification. To address these challenges, we introduce GAUDI (Graph Autoencoder Uncovering Descriptive Information), a novel unsupervised geometric deep learning framework designed to capture both local details and global structure. GAUDI employs an innovative hourglass architecture with hierarchical pooling and upsampling layers linked through skip connections, which preserve essential connectivity information throughout the encoding–decoding process. Even though identical or highly similar underlying parameters describing a system’s state can lead to significant variability in graph realizations, GAUDI consistently maps them into nearby regions of a structured and continuous latent space, effectively disentangling invariant process-level features from stochastic noise. We demonstrate GAUDI’s versatility across multiple applications, including small-world networks modeling, characterization of protein assemblies from super-resolution microscopy, analysis of collective motion in the Vicsek model, and identification of age-related changes in brain connectivity. Comparison with related approaches highlights GAUDI’s superior performance in analyzing complex graphs, providing new insights into emergent phenomena across diverse scientific domains.
@@ Introduction
Devising an unsupervised method capable of creating a structured and continuous latent space where different realizations of the state of a complex system are mapped into nearby points is a challenging task and, to the best of our knowledge, has not been tackled before. Such a method must account for the inherent randomness of these systems, where the same global parameters can produce different realizations, resulting in graphs that differ significantly in terms of both connectivity and features. Therefore, the method must reach a level of abstraction that enables pattern recognition at multiple scales and the association of combinations of patterns with a global organization. This cannot be achieved with graph contrastive learning since structural and semantic augmentations inherently change the state of the system, breaking the continuity and consistency required for a meaningful latent space representation.
@@ Results
For graphs, autoencoders have primarily been applied for downstream tasks such as link prediction, node classification, and node clustering [23, 24, 25]. These approaches often do not use pooling and upsampling. Instead, they encode structural information into the nodes and reconstruct the connectivity matrix from the node features, focusing on local features rather than capturing the global structure of the graph. This choice restricts the network’s ability to learn global information about the complex system. Even recent graph contrastive learning methods designed to preserve global structural and semantic patterns [15] are not able to assemble a meaningful latent space, even while effective at improving performance on downstream tasks.
To address these limitations, we propose a graph-convolutional variational autoencoder with an “hourglass” structure (Fig. 1b). At all levels of compression, the encoder explicitly sends the adjacency matrix and the cluster assignment matrix, which encodes the graph’s partitioning during pooling, to the decoder via bridging connections. These connections are similar to the skip connections used in the U-Net [26]. The presence of skip connections ensures that the connectivity information is preserved and lets GAUDI focus on the graph’s overall structure and high-level features rather than just the local connections. The “hourglass” architecture is crucial for increasing the perceptive field without inducing oversmoothing [27] and capturing multiscale information, resulting in a parametric representation that retains critical details at the bottleneck.
@@ Discussion
A defining feature of GAUDI is its use of skip connections to pass connectivity information between the encoder and decoder, effectively decoupling semantic from structural information in the learning process. This design allows the latent space to encode higher-order graph features, such as node attribute interactions, local neighborhood roles, and broader structural properties like clusters or communities. By focusing on these complementary aspects, GAUDI avoids redundancy and ensures robustness against variability introduced by stochastic processes. Unlike traditional graph autoencoders, which reconstruct adjacency matrices directly and risk capturing noise, GAUDI incorporates structural regularization to preserve process invariants. This ensures the latent space reflects essential graph topology while remaining robust to randomness.
@@ Methods
where, $\mathbf{V}_{\text{conv}}^{l-1}=z$ in the first decoder block. Subsequently, the upsampled node features undergo a graph convolution operation, as defined in Equation (3), using the binary adjacency matrix, $\mathbf{A}$, from the corresponding encoder level $l_{\text{enc}}$.
@@
GAUDI is trained to minimize the difference in node and edge features from the decoder output compared to the preprocessed graphs, using the mean absolute error. In addition to this loss, there is a weighted Kullback-Leibler divergence loss [47], and a loss for the MinCut pooling [28]. The total loss is a weighted sum of the losses, with the weights $\alpha$ for the node feature reconstruction loss, $\beta$ for the edge feature reconstruction loss, $\gamma$ for the Kullback-Leibler divergence, and $\delta$ for the MinCut pooling loss. The MinCut pooling is defined as
$L_{\text{MinCut}}=-\frac{Tr(S^{T}AS)}{Tr(S^{T}DS)}+\left\lVert\frac{\mathbf{S}^{T}\mathbf{S}}{\left\lVert\mathbf{S}^{T}\mathbf{S}\right\rVert_{F}}-\frac{\mathbf{I}_{\nu}}{\sqrt{\nu}}\right\rVert_{F}$ (8)
with $\left\lVert.\right\rVert_{F}$ denoting the Frobenius norm, $D$ denoting the degree matrix of $A$, and $\nu$ denoting the number of nodes of the pooled graph.
We use the following parameters to obtain the described results. All of the fully connected layers of GAUDI have a dimension of 96, unless specified otherwise. The latent space dimension is set to 8. The graphs are pooled down to $\nu_{1}=20$ nodes and $\nu_{2}=5$ nodes in the blocks of the encoder. The threshold for the binarization of the adjacency matrices after the pooling are set to $\tau_{1}=1/19$ and $\tau_{2}=1/5$. The choice of values for the thresholds are motivated by the aim to keep at least one connection for each node of the pooled graph, while not having a fully connected graph.

### v3

@@ Abstract
Graphs provide a powerful framework for modeling complex systems, but their structural variability poses significant challenges for analysis and classification. To address these challenges, we introduce GAUDI (Graph Autoencoder Uncovering Descriptive Information), a novel unsupervised geometric deep learning framework designed to capture both local details and global structure. GAUDI employs an hourglass architecture with hierarchical pooling and upsampling layers linked through skip connections, which preserve essential connectivity information throughout the encoding–decoding process. Even though identical or highly similar underlying parameters describing a system’s state can lead to significant variability in graph realizations, GAUDI consistently maps them into nearby regions of a structured and continuous latent space, effectively disentangling invariant process-level features from stochastic noise. We demonstrate GAUDI’s versatility across multiple applications, including small-world networks modeling, characterization of protein assemblies from super-resolution microscopy, analysis of collective motion in the Vicsek model, and identification of age-related changes in brain connectivity. Comparison with related approaches highlights GAUDI’s superior performance in analyzing complex graphs, providing new insights into emergent phenomena across diverse scientific domains.
@@ Introduction
Devising an unsupervised method capable of creating a structured and continuous latent space in which different realizations of the state of a complex system are mapped to nearby points is a challenging task and, to the best of our knowledge, has not been demonstrated in existing approaches. Such a method must account for the inherent randomness of these systems, where the same global parameters can produce different realizations, resulting in graphs that differ significantly in terms of both connectivity and features. Therefore, the method must reach a level of abstraction that enables pattern recognition at multiple scales and the association of combinations of patterns with a global organization. This cannot be achieved with graph contrastive learning since structural and semantic augmentations inherently change the state of the system, breaking the continuity and consistency required for a meaningful latent space representation.
@@ Results
For graphs, autoencoders have primarily been applied for downstream tasks such as link prediction, node classification, and node clustering [23, 24, 25]. These approaches often do not use pooling and upsampling. Instead, they encode structural information into the nodes and reconstruct the connectivity matrix from the node features, focusing on local features rather than capturing the global structure of the graph. This choice restricts the network’s ability to learn global information about the complex system. Even recent graph contrastive learning methods designed to preserve global structural and semantic patterns [15] often prioritize downstream task performance, without explicitly promoting the formation of a meaningful latent space.
To address these limitations, we propose a graph-convolutional variational autoencoder with an “hourglass” structure (Fig. 1b). At all levels of compression, the encoder explicitly sends the adjacency matrix and the cluster assignment matrix, which encodes the graph’s partitioning during pooling, to the decoder via bridging connections. These connections are similar to the skip connections used in the U-Net [26]. The skip connections help preserve connectivity information, reducing the need for coarser-scale representations to retain local details and thereby encourage the learning of representations that capture the graph’s overall structure and high-level features. The “hourglass” architecture is crucial for increasing the perceptive field without inducing oversmoothing [27] and capturing multiscale information, resulting in a parametric representation that retains critical details at the bottleneck.
@@ Discussion
A defining feature of GAUDI is its hierarchical encoder-decoder architecture, which applies graph convolutional layers at multiple scales with pooling operations that progressively reduce the size of the graph. This design enables information to be propagated both locally and globally across the graph while being compressed into a latent representation. Skip connections between the encoder and decoder complement this hierarchical design by preserving connectivity information. By integrating information across scales, GAUDI’s architecture reduces the reliance on local connectivity patterns alone and encourages the latent space to capture higher-order graph features, such as node attribute interactions, local neighborhood roles, and broader structural properties like clusters or communities. As a result, GAUDI learns compact graph representations that reflect essential graph topology while remaining robust to variability introduced by stochastic processes.
@@ Methods
where, $\mathbf{V}_{\text{dec}}^{l-1}=z$ in the first decoder block. Subsequently, the upsampled node features undergo a graph convolution operation, as defined in Equation (3), using the binary adjacency matrix, $\mathbf{A}$, from the corresponding encoder level $l_{\text{enc}}$.
@@
GAUDI is trained to minimize the difference in node and edge features from the decoder output compared to the preprocessed graphs, using the mean absolute error. In addition to this loss, there is a weighted Kullback-Leibler divergence loss [47], and a loss for the MinCut pooling [28]. The total loss is a weighted sum of the losses, with the weights $\alpha$ for the node feature reconstruction loss, $\beta$ for the edge feature reconstruction loss, $\gamma$ for the Kullback-Leibler divergence, and $\delta$ for the MinCut pooling loss. The MinCut pooling loss is defined as
$\mathcal{L}_{\text{MinCut}}=\sum_{l\in\mathcal{P}}\biggl(-\frac{Tr((\mathbf{S}^{l})^{T}\mathbf{A}^{l}\mathbf{S}^{l})}{Tr((\mathbf{S}^{l})^{T}\mathbf{D}^{l}\mathbf{S}^{l})}+\left\lVert\frac{(\mathbf{S}^{l})^{T}\mathbf{S}^{l}}{\left\lVert(\mathbf{S}^{l})^{T}\mathbf{S}^{l}\right\rVert_{F}}-\frac{\mathbf{I}_{N_{v}^{l+1}}}{\sqrt{N_{v}^{l+1}}}\right\rVert_{F}\biggr)$ (8)
with $\left\lVert.\right\rVert_{F}$ denoting the Frobenius norm, $\mathbf{D}^{l}$ denoting the degree matrix of $\mathbf{A}^{l}$, and $N_{v}^{l+1}$ denoting the number of nodes of the pooled graph. The set $\mathcal{P}$ indexes the encoder layers at which MinCut pooling is applied, including all encoder layers except the final one, where global pooling is performed.
We use the following parameters to obtain the described results. All of the fully connected layers of GAUDI have a dimension of 96, unless specified otherwise. The latent space dimension is set to 8. The graphs are pooled down to $N_{v}^{1}=20$ nodes and $N_{v}^{2}=5$ nodes in the blocks of the encoder. The threshold for the binarization of the adjacency matrices after the pooling are set to $\tau_{1}=1/19$ and $\tau_{2}=1/5$. The choice of values for the thresholds are motivated by the aim to keep at least one connection for each node of the pooled graph, while not having a fully connected graph.